a) Plot currents through the capacitor and the inductance of a series RLC circuit at resonance on
the same graph as functions of time. (b) What is the phase difference between the current and the
total voltage at frequency w>w0 , such that current amplitude is Imax/√2 ?
1
Expert's answer
2020-05-19T10:53:11-0400

a) The currents will be equal. Moreover, the magnitude of the current will be maximum:



b) Assume that we have voltage source V, resistance R, capacitance C, and inductance L. First, find the resonant frequency:


"\omega_0=\frac{1}{\sqrt{LC}}<\omega."


The maximum value of current is at resonance:


"I_\text{max}=\frac{V}R{}."

We need a current of


"I=\frac{I_\text{max}}{\sqrt2}=\frac{V}{R\sqrt2}."

Thus, the impedance of a series RLC circuit is


"Z=\sqrt{R^2+(X_L-X_C)^2}=R\sqrt2,"


and since the phase difference is


"\phi=\text{atan}\frac{X_L-X_C}{R},"

all we need is to find "(X_L-X_C)" from the equation above for impedance Z:


"X_L-X_C=R,
\\
\phi=\text{atan}\frac{R}{R}=0.785\text{ rad},"

or 45°.